What is the average kinetic energy (in kJ mol-1) of a hypothetical gas with molecular weight 12.8 g mol-1 at a temperature of 24.7 °C? Do not try to identify the gas.


1
Expert's answer
2022-02-09T15:40:03-0500

The average kinetic energy of a gas is determined by the following equation:

"E_T= \frac{3}{2} RT"

where "R" is the universal gas constant 8.314 "J mol^{-1} K^{-1}" J mol-1 K-1, "T" is a temperature. The average kinetic energy is:

"E_T= 8.314 \times (24.7+273) K = 3714.5 J/mol =3.71 kJ/mol"